Two motors are winding up rope. The first motor can wind up rope at 2ft per second, and the second motor can wind up rope at 5ft per second.
a). write a simplified expression that gives the length of rope by both motors in x seconds.
b). find the total length of rope wound up in 3 minutes.
c). how many minutes would it take to wind up 2100ft of rope by using both motors.
First off, i attempted to answer from the bottom up.
c). t=d/r i got the answer being 5 minutes.
b). d=rt 7ft per second (3min) 7ftps= 420ft per min answer would be 1260 ft
i'm just not quite sure how to write these answers correctly showing my work.

Answer by mananth(16946)

speed of DISABLED_event_one= 2 ft/s
speed of second=3ft/s
..
a. in x seconds they wind up together 2x+3x = 5x feet
b. 3 minutes = 180 second
in 3 minutes they will wind 5*180 = 900 ft
c. 2100 feet
together they wind 5 ft/s
so they will take 2100/5 = 420 seconds = 7 minutes
m.ananth@hotmail.ca